# A Closed-Form Solution for Kernel Adaptive Filtering

### v2

@@ 2.3 Nonlinear MMSE Solution
As before, let $X=\{X_{t},t\in T\}$ be a real-valued random process where each $X_{t}$ is a r.v. defined on a probability space $(\Omega,\mathcal{A},P)$, and $T$ is a compact subset of a separable metric space. Let $\mathcal{F}_{\mathcal{X}}=\{f_{x},x\in\mathcal{X}\}$ be a family of functions $f_{x}:\mathbb{R}\mapsto\mathbb{R}$ indexed by the elements of a compact set $\mathcal{X}$ such that $\mathbb{E}[|f_{x}(X_{t})|^{2}]<\infty$ for all $x\in\mathcal{X}$ and $t\in T$. From [26], we have $L_{2}(\Omega,\mathcal{A},P)$ as the Hilbert space of r.v.s in $(\Omega,\mathcal{A},P)$ with finite second-order moments. We can define the set $f(X)=\{f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T\}$ as a family of finite second-order random functions indexed by $x\in\mathcal{X}$ and $t\in T$. This set corresponds to the set of all r.v.s that can be written in the form:
$W=\sum\limits_{i=1}^{n_{W}}\sum\limits_{j=1}^{m_{W}}a_{ij}f_{x_{i}}(X_{t_{j}})$ (11)
for some positive integers $n_{W}$ and $m_{W}$. From the conditions defined above, we have that $W\in L^{2}(\Omega,\mathcal{A},P)$. By defining the inner product using the expected value we can form a linear manifold, $L(f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T)$ ,
$\begin{split}\langle W,V\rangle_{L^{2}}=\mathbb{E}[WV]&=\sum\limits_{i=1}^{n_{W}}\sum\limits_{j=1}^{m_{W}}\sum\limits_{k=1}^{n_{V}}\sum\limits_{\ell=1}^{m_{V}}a_{ij}b_{k\ell}\mathbb{E}[f_{x_{i}}(X_{t_{j}})f_{x_{k}}(X_{t_{\ell}})]\\ &=\sum\limits_{i=1}^{n_{W}}\sum\limits_{j=1}^{m_{W}}\sum\limits_{k=1}^{n_{V}}\sum\limits_{\ell=1}^{m_{V}}a_{ij}b_{k\ell}U(t_{j},t_{\ell},x_{i},x_{k})\end{split}$ (12)
@@
and by completing the space by adding the limits to all Cauchy sequences, we define the Hilbert space $L^{2}(f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T)$, or $L^{2}(f(X))$ for short. From (12) it is easy to see that for $s,t\in T$ and $x,y\in\mathcal{X}$ we can write the function $U(t,s,x,y)$ as $U((x,t),(y,s))$, that is $U:(\mathcal{X}\times T)\times(\mathcal{X}\times T)\mapsto\mathbb{R}$, then $U$ is positive semi-definite.
$0\leq\mathbb{E}[|W|^{2}]=\sum\limits_{i=1}^{n_{Y}}\sum\limits_{j=1}^{m_{Y}}\sum\limits_{k=1}^{n_{Y}}\sum\limits_{\ell=1}^{m_{Y}}a_{ij}a_{k\ell}U((x_{i},t_{j}),(x_{k},t_{\ell}))$ (13)
for any $\mathbf{a}\neq\mathbf{0}$.
Similar to the linear case, we can write any inner product in $L^{2}(f(X))$ as a function of the positive semi-definite function $U((x_{i},t_{j}),(x_{k},t_{\ell}))$ and vice versa. Therefore, we see that
$\langle W,V\rangle_{L^{2}}=\langle W^{{}^{\prime}},V^{{}^{\prime}}\rangle_{\mathcal{H}_{U}}$ (14)
where $W,V\in L^{2}(f_{x}(X_{t}),x,t\in\mathcal{X}\times T)$ and $W^{{}^{\prime}},V^{{}^{\prime}}\in\mathcal{H}_{U}$.
@@
Equation (14) suggests that $L^{2}(f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T)$ and $\mathcal{H}_{U}$ are congruent. Therefore any MMSE solution in $L^{2}(f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T)$ has an exact representation in $\mathcal{H}_{U}$, and Parzen’s solution in $\mathcal{H}_{U}$ is possible.
If we let $x,y\in\mathcal{X}\subseteq\mathbb{R}$ and $f_{x}=\kappa(\cdot,x)$ where $\kappa(\cdot,\cdot)$ is a reproducing kernel, then $U((x,t),(y,s))=\mathbb{E}[\kappa(X_{t},x)\kappa(X_{s},y)]$, and the finite second-order moment constraint mentioned above is fulfilled. Therefore, the MMSE solution in $\mathcal{H}_{U}$ is the orthogonal projection of some desired r.v. $Z$ into the space. This function will be a function defined over $\mathcal{X}\times T$. The orthogonal projection can be written as,
$\rho(y,s)=\mathbb{E}[Zf_{y}(X_{s})]=\mathbb{E}[Z\kappa(X_{s},y)]$ (15)
Finally, the MMSE solution in $\mathcal{H}_{U}$ is
$\begin{split}Z^{*}=\langle\rho,X\rangle_{\mathcal{H}_{U}}&=\sum_{s\in T}\sum_{t\in T}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{X}}f_{x}(X_{t})U((x,t),(y,s))^{\dagger}\rho(y,s)\\ &=\sum_{s\in T}\sum_{t\in T}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{X}}\kappa(X_{t},x)U((x,t),(y,s))^{\dagger}\rho(y,s)\end{split}$ (16)
We refer to this solution in the nonlinear data-dependent RKHS, $\mathcal{H}_{U}$, as the Functional Wiener Filter (FWF).
2.3.1 Mercer’s theorem and a simple congruence
For the case were we use positive definite kernel $\kappa$ to define $f_{x}:=\kappa(\cdot,x)$ with $x\in\mathcal{X}$, where $\mathcal{X}$ is a compact space (for instance a closed interval in $\mathbb{R}^{d}$), we can define a congruence based on Mercer’s theorem. First, note that Mercer’s theorem allows us to decompose the kernel as $\kappa(x,y)=\sum\limits_{m=1}^{N_{\mathcal{H}_{\kappa}}}\lambda_{m}\psi_{m}(x)\psi_{m}(y)$, were $N_{\mathcal{H}_{\kappa}}$ is either finite or a countable $\infty$. Then $L^{2}(f(X))$ is congruent with $L^{2}(\psi_{m}(X_{t}),(t,m)\in T\times\mathbb{N})$ and consequently also congruent with $\mathcal{H}_{\bm{U}}$ where $\bm{U}:(T\times\mathbb{N})\times(T\times\mathbb{N})\mapsto\mathbb{R}$ is a positive definite kernel defined as:
$\bm{U}((t,m),(s,n))=\mathbb{E}\left[\psi_{m}(X_{t})\psi_{n}(X_{s})\right]$ (17)
This additional congruent relationship can be used to change the domain of $U(s,t,x,y)$ where $x$ and $y$ are from a potentially uncountable set, to a countable domain. This is an important step in later sections where we give a method for computing the FWF from realizations.
@@ 2.3.2 Generalizing $\mathcal{H}_{U}$
where $g_{i}\in\operatorname{span}\left\{f_{x},x\in\mathcal{X}\right\}$. Using tap-delay embedding of the random process, we can generalize to functions of vectors in $\mathbb{R}^{D}$. For a random process $X$ and a given set of relative times $\left[\tau_{1},\tau_{2},\dots,\tau_{D-1}\right]$, we can define the tap-delay version $\bm{X}^{D}=\{\bm{X}^{D}_{t},t\in T\}$, such that $\bm{X}_{t}^{D}=[X_{t},X_{t-\tau_{1}},X_{t-\tau_{2}},\cdots,X_{t-\tau_{D-1}}]^{\top}$. For simplicity, we assume that $t-\tau_{i}\in T$ for all $t\in T$ and all $i=1,2,\dots,D-1$
We now proceed just as before by defining the linear manifold $L(f_{x}(\bm{X}_{t}^{D}),(x,t)\in\mathcal{X}\times T)$ as the span of this family, where now $f_{x}:\mathbb{R}^{D}\mapsto\mathbb{R}$. Any r.v. in this manifold can be written as,
$W=\sum_{i=1}^{n_{W}}\sum_{j=1}^{m_{W}}a_{ij}f_{x_{i}}(\bm{X}_{t_{j}}^{D})$ (19)
Similarly, we denote the completion of this linear manifold by $L^{2}(f_{x}(\bm{X}_{t}^{D}),(x,t)\in\mathcal{X}\times T)$ or simply as $L^{2}(f(\bm{X}^{D}))$. Furthermore, we can extend the above notation to build nested sets. For example, if $T$ is the set of integers, we can choose a positive integer $L>0$ to build the nested set $\bm{X}_{t}^{DL}=\{\bm{X}_{t}^{D},\bm{X}_{t-1}^{D}...,\bm{X}_{t-(L-1)}^{D}\}$. We will refer to $D$ as the sample embedding size, and $L$ as the window size. By varying $D$ and $L$, the combinations of r.v.s in $X$ over which the functions in $\mathcal{H}_{U}$ are defined can be adjusted.
@@
These nested sets will give the building blocks for creating a spectrum of spaces with different levels of generality and computation requirements. On one end of this spectrum are less general spaces that will require less computation to compute the MMSE solution and less data for convergence to this MMSE solution. On the other end of this spectrum are more general spaces, which require more computation to compute the MMSE solution and often more data to converge to the MMSE solution. These trade-offs between error, computational complexity, and the necessary amount of training data for convergence are unavoidable and must be balanced via proper selection of the hyperparameters $L$ and $D$. Similar assumptions to those made above are commonplace in KAF, where an “embedding size” is selected that fixes the dimensionality of the inputs [17].
@@ 3 Computing the MMSE Solution
The sections above show theoretically how to extend Parzen’s idea of a MMSE in a data-dependent universal RKHS. Now we show one method for practically computing the nonlinear MMSE solution. We will begin by defining an explicit finite-dimensional RKHS that approximates the Gaussian RKHS, and approximate there the solution given in section 2.3.
@@ 3.1 The Explicit Finite-Dimensional Approximation for the Gaussian RKHS
In [14], an explicit mapping function based on the Taylor expansion of the Gaussian kernel is used to give a finite-dimensional approximation of the feature space specified by the Gaussian kernel. A full derivation of this explicit mapping function can be found in [3]. This is just one way of creating an explicit feature space; another notable technique employs Random Fourier features (RFF) [28], but it has higher computational costs than the Taylor expansion-based method. We will refer to this explicit finite-dimensional Hilbert Space as $\mathcal{H}_{S}$, with kernel $S(\cdot,\cdot):\mathbb{R}^{D}\times\mathbb{R}^{D}\rightarrow\mathbb{R}$. The explicit mapping function given in [3] (also see [41]) is written as
$\phi_{k,j}(x)=e^{-\frac{\|\mathbf{x}\|^{2}}{2\sigma^{2}}}\frac{1}{\sigma^{k}\sqrt{k!}}\prod_{i=1}^{k}x_{j_{i}}$ (20)
@@
The finite rank approximation of the Gaussian kernel, that is obtained by truncating the Taylor series expansion to the first $K$ terms,
@@
$=$ $e^{-\frac{\|x\|^{2}}{2\sigma^{2}}}e^{-\frac{\|x^{\prime}\|^{2}}{2\sigma^{2}}}\sum\limits_{k=0}^{K}\frac{\left(x^{\top}x^{\prime}\right)^{k}}{\sigma^{2k}k!}=\sum_{m=1}^{M}\phi_{m}(x)\phi_{m}(x^{\prime})$
@@
Since the Gaussian kernel over a closed bounded interval is a Mercer kernel, the representations based on the full Taylor expansion and the eigendecomposition of the integral operator induced by the kernel (Mercer’s theorem) are equivalent and thus the closure of the span of $\{\phi_{m}(X_{t}),(m,t)\in\mathbb{N}\times T\}$ contains $L^{2}(f(X))$. The truncated series yields a family of functions $\{\phi_{m}(X_{t}),(m,t)\in[1,M]\times T\}$ where we can approximate any $f_{x}=G_{\sigma}(\cdot,x)$ by a finite superposition:
$f_{x}(x^{\prime})\approx\sum\limits_{m=1}^{M}\phi_{m}(x)\phi_{m}(x^{\prime}),$ (22)
Due to the exponential decay of the Gaussian kernel’s eigenvalues, effective finite-rank approximations can often be obtained with just a few features in the expansion [14]. The main advantage of the explicit feature space in the context of this work is that it simplifies the design of linear operators in the feature space because we can represent them as finite-dimensional matrices. Also, since operators can be represented as matrices and functionals as vectors, computations in $\mathcal{H}_{S}$ enjoy the computational tractability of matrix and vector multiplications.
We should draw attention to some deficiencies of this finite-dimensional RKHS. In [34], an overview of universal and characteristic reproducing kernels is given. A reproducing kernel is c-universal if any target function can be approximated to an arbitrary degree of accuracy, provided the search algorithm converges to the desired function and the number of training samples goes to infinity. A reproducing kernel is called characteristic if the kernel mean embedding is an injective function. In [22] it is shown that a kernel is c-universal if and only if it is characteristic. The Gaussian kernel is both c-universal and characteristic. However, its finite approximation, $S(\cdot,\cdot)$, loses these properties. As an example, consider two probability measures say $p(x)$ and $q(x)$ whose kernel mean embeddings using the Gaussian kernel have feature space representations, $\{\phi_{n}(p(x))\}^{\infty}_{n=1}$ and $\{\phi_{n}(q(x))\}^{\infty}_{n=1}$. There may be pairs of probability measures such that
$\phi_{n}(p(x))=\phi_{n}(q(x))\quad\forall n\leq k,\quad\phi_{n}(p(x))\neq\phi_{n}(q(x))\quad\forall n>k$ (23)
Since $S(\cdot,\cdot)$ is a truncation of this feature space, two probability measures satisfying (23) cannot be disambiguated in $\mathcal{H}_{S}$ and therefore the mean embedding in $\mathcal{H}_{S}$ is not injective. This also means that $\mathcal{H}_{S}$ is not universal. While these issues are theoretically consequential, due to the eigenvalue roll-off of the Gaussian kernel, we expect that the effects of this deficiency of $\mathcal{H}_{S}$ will be manageable. This assumption will be experimentally validated in later sections.
In previous works such as [14], [28], and [41], an explicit mapping function is used to decouple model size from the number of training samples. While our method inherits this as a strength of using an explicit approximation, we will also see that it is necessary to work with individual features of the RKHS if we want to define a closed-form MMSE solution that is practically computable. Therefore, the explicit RKHS provides an elegant and efficient way of performing the operations necessary for our closed-form solution.

### v3

@@ 2.2 Nonlinear MMSE Solution
As before, let $X$ be a real-valued random process where each $X_{t}$ is a r.v. defined on a probability space $(\Omega,\mathcal{A},P)$, and $T$ is a compact subset of a separable metric space. Let $\mathcal{F}_{\mathcal{X}}=\{f_{x},x\in\mathcal{X}\}$ be a family of functions $f_{x}:\mathbb{R}\mapsto\mathbb{R}$ indexed by the elements of a compact set $\mathcal{X}$ such that $\mathbb{E}[|f_{x}(X_{t})|^{2}]<\infty$ for all $x\in\mathcal{X}$ and $t\in T$. Note that if we let $x,y\in\mathcal{X}\subseteq\mathbb{R}$ and $f_{x}=\kappa(\cdot,x)$ where $\kappa(\cdot,\cdot)$ is a reproducing kernel, then these constraints will be met. From [27], we have $L_{2}(\Omega,\mathcal{A},P)$ as the Hilbert space of r.v.s. in $(\Omega,\mathcal{A},P)$ with finite second-order moments. We can define the set $f(X)=\{f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T\}$ as a family of finite second-order random functions indexed by $x\in\mathcal{X}$ and $t\in T$. This set corresponds to the set of all r.v.s. that can be written in the form, $W=\sum\limits_{i=1}^{n_{W}}\sum\limits_{j=1}^{m_{W}}a_{ij}f_{x_{i}}(X_{t_{j}})$ for some positive integers $n_{W}$ and $m_{W}$. From the conditions defined above, we have that $W\in L^{2}(\Omega,\mathcal{A},P)$. By defining the inner product using the expected value as,
$\begin{split}\langle W,V\rangle_{L^{2}}=\mathbb{E}[WV]&=\sum\limits_{i=1}^{n_{W}}\sum\limits_{j=1}^{m_{W}}\sum\limits_{k=1}^{n_{V}}\sum\limits_{\ell=1}^{m_{V}}a_{ij}b_{k\ell}\mathbb{E}[f_{x_{i}}(X_{t_{j}})f_{x_{k}}(X_{t_{\ell}})]\\ &=\sum\limits_{i=1}^{n_{W}}\sum\limits_{j=1}^{m_{W}}\sum\limits_{k=1}^{n_{V}}\sum\limits_{\ell=1}^{m_{V}}a_{ij}b_{k\ell}U(t_{j},t_{\ell},x_{i},x_{k}),\end{split}$ (4)
@@
we can form a linear manifold, $L(f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T)$. Then, by adding the limits to all Cauchy sequences, we define the Hilbert space $L^{2}(f_{x}(X_{t}),(x,t)\in\mathcal{X}\times T)$, abbreviated as $L^{2}(f(X))$. From (4) it is easy to see that for $s,t\in T$ and $x,y\in\mathcal{X}$ we can write the function $U(t,s,x,y)$ as $U((x,t),(y,s))$. This operator, $U:(\mathcal{X}\times T)\times(\mathcal{X}\times T)\mapsto\mathbb{R}$, is a positive semi-definite function, and is therefore the kernel for some RKHS, $\mathcal{H}_{U}$.
@@
Equation (4) suggests that $L^{2}(f(X))$ and $\mathcal{H}_{U}$ are congruent. Therefore, any MMSE solution in $L^{2}(f(X))$ has an exact representation in $\mathcal{H}_{U}$, and Parzen’s solution in $\mathcal{H}_{U}$ is possible. The MMSE solution in $\mathcal{H}_{U}$ is the orthogonal projection of some desired r.v. $Z$ into the space. This projection can be written as $\rho(y,s)=\mathbb{E}[Zf_{x}(X_{s})]$. Finally, the MMSE solution in $\mathcal{H}_{U}$, which we call the Functional Wiener Filter (FWF), is given as
$Z^{*}=\langle\rho,X\rangle_{\mathcal{H}_{U}}=\sum_{s\in T}\sum_{t\in T}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{X}}f_{x}(X_{t})U((x,t),(y,s))^{\dagger}\rho(y,s).$ (5)
@@ 2.2.1 Generalizing $\mathcal{H}_{U}$
where $g_{i}\in\operatorname{span}\left\{f_{x},x\in\mathcal{X}\right\}$. This is restrictive if we wish to construct nonlinear functions of more than one sample. Using tap-delay embedding of the random process, we can generalize (6) to functions of vectors in $\mathbb{R}^{D}$. For a random process $X$ and a given set of relative times $\left[\tau_{1},\tau_{2},\dots,\tau_{D-1}\right]$, we can define the tap-delay version $\bm{X}^{D}=\{\bm{X}^{D}_{t},t\in T\}$, where $\bm{X}_{t}^{D}=[X_{t},X_{t-\tau_{1}},X_{t-\tau_{2}},\cdots,X_{t-\tau_{D-1}}]^{\top}$. For simplicity, we assume that $t-\tau_{i}\in T$ for all $t\in T$ and all $i=1,2,\dots,D-1$.
We now proceed just as before by defining the linear manifold $L(f_{x}(\bm{X}_{t}^{D}),(x,t)\in\mathcal{X}\times T)$ as the span of this family, where now $f_{x}:\mathbb{R}^{D}\mapsto\mathbb{R}$. Any r.v. in this manifold can be written as $W=\sum_{i=1}^{n_{W}}\sum_{j=1}^{m_{W}}a_{ij}f_{x_{i}}(\bm{X}_{t_{j}}^{D})$. Similarly, we denote the completion of this linear manifold by $L^{2}(f_{x}(\bm{X}_{t}^{D}),(x,t)\in\mathcal{X}\times T)$ abbreviated as $L^{2}(f(\bm{X}^{D}))$. Furthermore, we can extend the above notation to build nested sets. For example, if $T$ is the set of integers, we can choose a positive integer $L>0$ to build the nested set $\bm{X}_{t}^{DL}=\{\bm{X}_{t}^{D},\bm{X}_{t-1}^{D}...,\bm{X}_{t-(L-1)}^{D}\}$ and $\bm{X}^{DL}=\{\bm{X}_{t}^{DL},t\in T\}$. We will refer to $D$ as the sample embedding size, and $L$ as the window size. By varying $D$ and $L$, the combinations of r.v.s. in $X$ over which the functions in $\mathcal{H}_{U}$ are defined can be adjusted. Figure 1 gives a visual depiction of the sample embedding scheme.
@@
In later sections, these nested sets will give the building blocks for creating a spectrum of spaces with different levels of generality and computation requirements. An overview of the trade-offs controlled by the hyperparameters of the FWF is given in section 3.5.
@@ 3 Computing the MMSE Solution
The sections above show theoretically how to extend Parzen’s idea of a MMSE in a data-dependent and potentially universal RKHS. Now, we give one method for practically computing this nonlinear MMSE solution. We first define a congruence which shifts the domain of $U(s,t,x,y)$ where $x$ and $y$ are from a potentially uncountable set, to a countable domain. We then introduce an explicit finite-dimensional RKHS that approximates the Gaussian RKHS. Finally, we approximate the solution given in section 2.2 in this finite-dimensional RKHS.
3.0.1 Mercer’s theorem and a simple congruence
For the case where we use a positive definite kernel $\kappa$ to define $f_{x}:=\kappa(\cdot,x)$ with $x\in\mathcal{X}$, where $\mathcal{X}$ is a compact space (for instance a closed interval in $\mathbb{R}^{d}$), we can define a congruence based on Mercer’s theorem. This congruence can be used to change the domain of $U(s,t,x,y)$ where $x$ and $y$ are from a potentially uncountable set, to a countable domain. First, note that Mercer’s theorem allows us to decompose the kernel as $\kappa(x,y)=\sum\limits_{m=1}^{N_{\mathcal{H}_{\kappa}}}\lambda_{m}\psi_{m}(x)\psi_{m}(y)$, where $N_{\mathcal{H}_{\kappa}}$ is either finite or countably infinite. Then $L^{2}(\kappa(X))$ is congruent with $L^{2}(\psi_{m}(X_{t}),(t,m)\in T\times\mathbb{N})$ and consequently also congruent with $\mathcal{H}_{\bm{U}}$ where $\bm{U}:(T\times\mathbb{N})\times(T\times\mathbb{N})\mapsto\mathbb{R}$ is a positive definite kernel defined as,
$\bm{U}((t,m),(s,n))=\mathbb{E}\left[\psi_{m}(X_{t})\psi_{n}(X_{s})\right],$ (7)
where $\bm{U}(s,t,n,m)$ is now defined only over countable sets.
@@ 3.1 The Explicit Finite-Dimensional Approximation for the Gaussian RKHS
In [8], an explicit mapping function based on the Taylor expansion of the Gaussian kernel ($G(\cdot,\cdot)$) is used to give a finite-dimensional approximation of the feature space specified by the Gaussian kernel. A full derivation of this explicit mapping function can be found in [29]. This is just one way of creating an explicit feature space; another notable technique employs Random Fourier features (RFF) [30], but it has higher computational costs than the Taylor expansion-based method. We will refer to this explicit finite-dimensional Hilbert Space as $\mathcal{H}_{S}$, with kernel $S(\cdot,\cdot):\mathbb{R}^{D}\times\mathbb{R}^{D}\rightarrow\mathbb{R}$. The explicit mapping function given in [29] (also see [31]) is written as
$\phi_{k,j}(x)=e^{-\frac{\|\mathbf{x}\|^{2}}{2\sigma^{2}}}\frac{1}{\sigma^{k}\sqrt{k!}}\prod_{i=1}^{k}x_{j_{i}},$ (8)
@@
The finite rank approximation of the Gaussian kernel is then obtained by truncating the Taylor series expansion to the first $K$ terms,
@@
$=$ $e^{-\frac{\|x\|^{2}}{2\sigma^{2}}}e^{-\frac{\|x^{\prime}\|^{2}}{2\sigma^{2}}}\sum\limits_{k=0}^{K}\frac{\left(x^{\top}x^{\prime}\right)^{k}}{\sigma^{2k}k!}=\sum_{m=1}^{M}\phi_{m}(x)\phi_{m}(x^{\prime}).$
@@
Since the Gaussian kernel over a closed bounded interval is a Mercer kernel, the representations based on the full Taylor expansion and the eigendecomposition of the integral operator induced by the kernel (Mercer’s theorem) are equivalent. Therefore, the congruent relationship detailed in 3.0.1 applies to the Taylor expansion, and the closure of the span of $\{\phi_{m}(X_{t}),(m,t)\in\mathbb{N}\times T\}$ contains $L^{2}(\{G_{x}(X_{t}),(x,t)\in\mathcal{X}\times T\})$. Truncating this series yields a family of functions, $\{\phi_{m}(X_{t}),(m,t)\in[1,M]\times T\}$, where we can approximate any $f_{x}=G(\cdot,x)$ by a finite superposition, $f_{x}(x^{\prime})\approx\sum\limits_{m=1}^{M}\phi_{m}(x)\phi_{m}(x^{\prime})$.
In previous works [8, 30, 31] an explicit mapping function is used to decouple model size from the number of training samples. While our method inherits this as a strength of using an explicit approximation, the main advantage of the explicit feature space in the context of this work is that it simplifies the design of linear operators because we can represent them as finite-dimensional matrices. This allows for a practical method for computing the closed-form solution detailed in 2.2. The drawback of the finite-dimensional RKHS is that it is no longer universal. However, it is shown in [8], and in later sections (also see supplementary materials), that effective finite-rank approximations can be obtained with just a few features in the expansion.